# MLC at Feigenbaum points

### v2

@@ 6.1.3. The spiraling case:
The argument of this section will be repeated in §7.0.1 with ${\mathfrak{B}}^{[1]}_{a+1}$ being replaced by a little Julia set ${\mathfrak{K}}^{[1]}_{0}$. Observe that no specific properties of ${\mathfrak{B}}^{[1]}_{a+1}$ has been used in this section (only that it is a forward invariant set that is disjoint from ${\mathfrak{B}}^{[1]}$).
@@ 7. Conclusions
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\leq K_{\overline{p}}{\ \ }{\ \ }{\ \ }\text{ for }{\ \ }m\ \gg_{{\mathcal{W}}(F)}\ 1,$
@@
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\succeq_{F}C^{nm}_{\delta}K.$
@@
(7.1) ${\mathcal{W}}_{\bullet}(G_{m})\leq K_{\overline{p}}\hskip 28.45274pt\text{ for}m\gg_{{\mathcal{W}}(F)}1$
@@
$K\coloneqq{\mathcal{W}}\left({\mathcal{G}}\right)\gg_{K_{\overline{p}}}1,\hskip 19.91684pt\text{ where }{\ \ }{\mathcal{G}}\coloneqq{\mathcal{F}}\left(V\setminus\bigcup_{i\not=0}{\mathfrak{B}}^{[1]}_{i},\ {\mathfrak{K}}^{[1]}_{0}\right).$

### v3

@@ 6.1.3. The spiraling case:
The argument of this section will be repeated in §7.0.1 with ${\mathfrak{B}}^{[1]}_{a+1}$ being replaced by a little Julia set ${\mathfrak{K}}^{[1]}_{0}$. Observe that no specific properties of ${\mathfrak{B}}^{[1]}_{a+1}$ has been used in this section (only that it is a forward invariant set that is disjoint from ${\mathfrak{B}}^{[1]}_{a}$).
@@ 7. Conclusions
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\leq K_{\overline{p}}{\ \ }{\ \ }{\ \ }\text{ for }{\ \ }m\ \gg_{{\mathcal{W}}_{\bullet}(F)}\ 1,$
@@
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\succeq_{F}\ C^{n(m-1)}_{\delta}K.$
@@
(7.1) ${\mathcal{W}}_{\bullet}(G_{m})\leq K_{\overline{p}}\hskip 28.45274pt\text{ for }m\gg_{{\mathcal{W}}_{\bullet}(F)}1$
@@
[…]